OpenStudy (jenniferjuice):

math question reward is 6 medals c:

OpenStudy (jenniferjuice):

What is the axis of symmetry for f(x) = −5x2 − 20x − 10? x = 6 x = −6 x = 2 x = −2

OpenStudy (anonymous):

first, factor out -5 = -5(x^2 +4x+2) = -5(x+2)(x+2) = -5(x+2)^2 refer to equation: = (x-h)^2 + k h= -2 Axis of symmetry is h. x= h = -2